OpenStudy (anonymous):

two perpendicular chords divide a circle with a radiusof 13 cm. into four parts, if the perpendicular distances of both chords are 5 cm. each from the center of the circle , find the area of the smallest part

OpenStudy (anonymous):

From the diagram, triangle AOG is right-angled at G, so AG = √(13^2 - 5^2) = 12 cm. Therefore, x = AC = BC = 12 - 5 = 7 cm. Therefore, AB = 7√2 and AD = BD = 7√2/2. Let θ = angle AOB. Then angle OAB = 90 - θ/2. In triangle AOB, (7√2)/sin(θ) = 13/sin(90 - θ/2). Therefore, θ = 2*arcsin(7√2/26). Area of sector AOB = (1/2)*13^2*2*arcsin(7√2/26) = 66.01259 cm^2. By Pythagoras, OC = 5√2, so area of triangle OCB = (1/2)*5√2*13*(7√2/26) = 35/2. If it's not clear, the multiplier, 7√2/26, is the value of sin(θ/2) (for angle DOB). Area of triangle OCA = area of triangle OCB, so 2*(area of triangle AOC) = 35 cm^2. Thus, grey area = Area of sector AOB - 2*(area of triangle AOC) = 66.01259 - 35 = 31.01259 cm^2 General formula (R = radius of circle, d = distance of both chords from centre) : Let k = √(R^2 - d^2) - d, and provided that d ≤ R/√2, then - Area of smallest part is A = R^2*arcsin[k/(R√2)] - dk In both k and A, replacing (d) with (-d) gives A = area of largest part. The other two equal areas are then easily obtained by subtraction from πR^2

hartnn (hartnn):

draw a figure first.

OpenStudy (anonymous):



OpenStudy (anonymous):

the wau that sir ??

hartnn (hartnn):

 did you get all those ??? i used the pythagoras theorem once